Circular No. 8316

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004bd IN NGC 3786 M. Armstrong, Rolvenden, U.K., reports his discovery of an apparent

supernova (mag ∼ 14.3) on unfiltered CCD images taken on Apr. 7.120 and 7.847 UT with a 0.35-m reflector. SN 2004bd is located at α = 11^{h}39^{m}42^{s}.18, δ = +31^{o}54′31′′.8 (equinox 2000.0), which is ≈ 4′′.7 west and 1′′.2 south of the center of NGC 3786. Nothing was visible at this position on Armstrong’s images taken on 2000 Dec. 30, 2003 May 25, and Dec. 7 or on Palomar Sky survey red and blue plates. M. Mobberley, Cockfield, U.K., confirmed the new object on an unfiltered CCD image taken on Apr. 7.831.

1999 DJ_{4}

P. Pravec, P. Kušnirák, and L. Šarounová, Ondřejov Observatory; P. Brown and N. Kaiser, University of Western Ontario (UWO); G. Masi, University of Rome ‘Tor Vergata’; and F. Mallia, Latina, Italy, report that photometric observations — obtained during Feb. 21–Mar. 28 at Ondřejov, UWO’s Elginfield Observatory, and the Tenagra II telescope near Nogales, AZ — show that the Apollo-type minor planet 1999 DJ_{4} is a binary system with an orbital period of 17.72 hr. The primary rotates with a period of 2.514 hr, and its lightcurve amplitude of 0.10 mag is indicative of a nearly spheroidal shape. Mutual eclipse/occultation events that are 0.2 mag deep indicate a secondary-to-primary mean-diameter ratio of 0.4 or greater. Further observations, especially radar and spectroscopic, are desirable.

V5114 SAGITTARII M. Soma, National Astronomical Observatory of Japan, reports that

V5114 Sgr will be occulted by the moon numerous times this year, providing the following dates and approximate areas of visibility: Apr. 10, China; May 8, western Europe; June 4, northern Japan; July 2, Europe; July 29, Japan; Aug. 25, Europe and northern Africa; Sept. 22, eastern USA.

Visual magnitude estimates: Mar. 17.190 UT, 8.0 (J. D. Shanklin, Rothera, Antarctica); 18.162, 9.4 (M. Lehký, Hradec Králové, Czech Rep.); 18.801, 8.8 (A. Pearce, Nedlands, W. Australia); 19.828, 8.9 (Pearce); 21.826, 9.1 (Pearce); 22.249, 9.3 (R. Y. Shida, São Paulo, Brazil). 23.865, 9.6 (Pearce, Noble Falls, W. Australia); 25.13, 9.5 (Shanklin, Stanley, Falkland Islands); 25.835, 9.8 (Pearce); 28.177, 10.0 (P. Schmeer, Bischmisheim, Germany); 29.172, 10.3 (Schmeer); 30.164, 10.4 (Schmeer); 30.459, 10.2 (J. D. West, Mulvane, KS); Apr. 2.778, 10.6 (S. Yoshida, Ibaraki, Japan).

2004 April 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green